Bank of Delta has $120 in reserves, $520 in loans and checkable deposits of $640. Assume the required reserve ratio is 9%. A new customer deposits $110 in a checking account. After the bank receives this new deposit, the bank transforms all its excess reserves into loans and consequently, the bank has total loans of?

$682.5
$633
$605.5
$595
Answer #1

Answer


total deposits =old deposits + new depsoits
=640+110
=750
required reserves =deposit *required reserves
=750*0.09
=67.5
loan amount =deposits -required reserves
=750-67.5
=682.5
the total laon with the bank is $682.5

option 1